OpenStudy (anonymous):

(x-8/7)(x+7/8). please help. Im lost after i re-write the second part

OpenStudy (anonymous):

\[\left(x-\frac{8}{7}\right) \left(x+\frac{7}{8}\right)=x^2-\frac{15 x}{56}-1 \]Mulitply by 56 and factor.\[56 x^2-15 x-56=(7 x-8) (8 x+7) \]Divided by 56.\[\frac{1}{56}(7 x-8) (8 x+7) \]